Domenik has a coin bank in which he saves only P1 coin and 5 coins. When the coin bank was full, he asked his father to open it for him and asked to count how much money was in it. His father told him that all in all Domenik was able to save p238 and there were 66 coins all in all. How many P1 and 5 coins where there in the coin bank?
Answer by mananth(16946)

Domenik has a coin bank in which he saves only P1 coin and 5 coins. When the coin bank was full, he asked his father to open it for him and asked to count how much money was in it. His father told him that all in all Domenik was able to save p238 and there were 66 coins all in all. How many P1 and 5 coins where there in the coin bank?
..
P1 = x numbers
P5 = 66-x numbers
..
Total value = P238
...
1x+5(66-x)=238
x+330-5x=238
-4x=238-330
-4x=-92
/-4
x=23 P1 coins
...
66-23 = 43 P5 coins
...
m.ananth@hotmail.ca